OpenStudy (anonymous):

Find the simplified form of the expression. Give your answer in scientific notation. (3x10^-6)(6x10^-8)

OpenStudy (anonymous):

Have you tried it??

OpenStudy (anonymous):

I did a little bit but then after a while it got confusing..

OpenStudy (unklerhaukus):

\[(3\times10^{-6})(6\times10{^{-8}})\]\[=(3\times6)(10^{-6}\times10{^{-8}})\]\[=(3\times6)(10^{-6-8})\]

OpenStudy (unklerhaukus):

note that a number in scientific notation will have only one digit before the decimal place,

OpenStudy (anonymous):

okay so um, is this the answer 18x10^-15

OpenStudy (anonymous):

no that is not scientific notation because \(18>10\)

OpenStudy (anonymous):

you get \[18\times 10^{-14}=1.8\times 10^{-13}\]

OpenStudy (anonymous):

why does the ^-14 turn into ^-13? can you explain how you did that?

OpenStudy (unklerhaukus):

\[18\times 10^{-14}\]\[=18\times 10^{-14}\times\frac{10}{10}\]\[=\frac{18}{10}\times 10^{-14+1}\]\[=1.8\times 10^{-13}\]

OpenStudy (anonymous):

okay, i think i kind of got it. thank you both!

OpenStudy (anonymous):

can one you check my answer to this one. (9*10^5)(6*10^-7) (9*6)(10^5^-7) 9*6=54=5.4 10^5^-7=10^-34 so my answer is 5.4x10^-34

OpenStudy (unklerhaukus):

\[(9\times10^5)(6\times10^{-7})\]\[=(9\times6)(10^5\times10^{-7})\]\[=(9\times6)(10^{5-7})\]

OpenStudy (anonymous):

if you ever get confused about which way to move the decimal if you get number larger than 10 use small numbers as an example by which i mean, suppose you had \(18\times 10^{-2}=18\times .01=.18\) and you wanted it in scientific notation you would need \(1.8\times 10^{-1}=.18\) so you see which direction to change the exponent

OpenStudy (anonymous):

k, thank you.

OpenStudy (anonymous):

(9*10^5)(6*10^?)

OpenStudy (anonymous):

i need help with (9*10^5)(6*10^?)